OpenStudy (anonymous):

find the coordinates of the turning points of y=2x^2(3-x) and determine their nature.

OpenStudy (cwrw238):

find the derivative and equate to zero y=2x^2(3-x) y = 6x^2 - 2x^3 do you know how to differentiate this?

OpenStudy (cwrw238):

use the formula if y = ax^n y' = anx^(n-1) so for example if y = 3x^4 y' = 3*4 x^(4-1) = 12 x^3